# Towards Agentic AI on Particle Accelerators

### v1

@@ Abstract
As particle accelerators grow in complexity, traditional control methods face increasing challenges in achieving optimal performance. This paper envisions a paradigm shift: a decentralized multi-agent framework for accelerator control, powered by Large Language Models (LLMs) and distributed among autonomous agents. We present a proposition of a self-improving decentralized system where intelligent agents handle high-level tasks and communication and each agent is specialized control individual accelerator components.
This approach raises some questions: What are the future applications of AI in particle accelerators? How can we implement an autonomous complex system such as a particle accelerator where agents gradually improve through experience and human feedback? What are the implications of integrating a human-in-the-loop component for labeling operational data and providing expert guidance? We show two examples, where we demonstrate viability of such architecture.
@@ 1 Introduction
Particle accelerator operation involves a complex interplay of interrelated systems, each requiring precise control and optimization. Traditionally, these systems have been managed through a combination of human expertise and highly specialized algorithms with recent developments including machine learning techniques [7] where some are successfully deployed in operation like e.g[11, 8, 29].
Most of these algorithms are designed to operate in a narrow domain of accelerator functionality. The systems have noise, tight operational limits, and deterministic interplay through the control system to ensure human and equipment safety, preventing scenarios that could lead to hazardous conditions or equipment damage. While effective, this approach can sometimes struggle to achieve optimal overall performance due to the challenges of integrating these disparate systems. Even when a sufficiently comprehensive model of the facility is constructed, it can be highly sensitive and prone to problems as the accelerator transitions through different states or experiences drifts and thus requires human intervention for setup and ongoing maintenance.
This paper proposes a paradigm shift in accelerator control systems, envisioning a future where self-improving agents control separated sub-components using algorithms that inform operators of significant events. In this framework, we show how autonomous agents controlled by LLMs serve not only in the creation and retrieval of accelerator operation documentation but also as high-level coordinators, handling complex tasks such as the interpretation of the machine state from analyzing control system channels or inter-agent communication with a degree of autonomy in decision-making and communication. Meanwhile, dedicated agents controlled by LLMs execute specific, well-established calculations or operations within their domains of expertise. This decentralized approach, illustrated in our conceptual diagram Fig. 1, showcases a potential future for the operation of particle accelerators with the flexibility to deploy more advanced agents.
@@
Figure 1: Schematic overview of a decentralized agent-based control architecture. The left diagram illustrates the modular structure where each agent, consisting of devices, algorithms, and agent controllers, manages specific subsystems. The right figure shows these agents interact with both the physical components of the accelerator and the control room, enabling prompting current state, monitoring, decision-making, and adaptive responses to complex operational scenarios.
@@ 2 Related Work
Machine learning techniques have increasingly been applied to various aspects of accelerator physics in recent years. Edelen et al[7] provide a comprehensive review of machine learning applications in particle accelerators, covering areas such as beam diagnostics, control systems, and performance optimization.
@@ 2.1 Agentic AI and LLMs in Complex Reasoning Tasks
There is a rise of agentic AI that has demonstrated remarkable capabilities in complex reasoning tasks [2]. According to [27], an effective agent in an agentic system should possess autonomy to operate independently, reactivity to respond to environmental changes, and proactiveness to take initiative in pursuing its objectives, enabling it to function intelligently and adaptively in complex environments. We will highlight the most relevant works that align with our problem.
@@
The field of use of AI agents to effectively engage with their environment and complete a wide array of tasks has been gaining more and more popularity. Park et al[19] introduce generative agents that use LLMs to simulate human behavior in interactive environments, showcasing how LLMs can be utilized to model intricate multi-agent interactions and environmental complexities. Wang et al[25] uses an LLM-powered agent to autonomously explore, acquire skills, and make discoveries in unfamiliar environments using components like an automatic curriculum and skill library. AgentVerse [3] shows a framework for collaborative groups of LLM-powered expert agents, showing improved performance over single agents in various tasks. This multi-agent approach leverages LLMs for high-level tasks while enabling specialized components to work together autonomously on complex problems. AutoGen [26] framework allows building LLM applications using multiple conversable agents. They show that autonomous, communicating agents can control subcomponents of a system. Particle accelerator control often utilizes operation sequencing [9, 1], there LLM+P [17] can help by translating natural language into formal planning languages, potentially bridging the gap between planning, natural language, and accelerator controls. To deal with environment changes, Shinn et al[23] proposes Reflexion, a framework that reinforces language agents through linguistic feedback rather than traditional weight updates. Reflexion agents verbally reflect on task feedback and maintain reflective text in an episodic memory buffer to improve decision-making in subsequent trials. To impose the factuality, [6] proposes multiagent debate where multiple LLMs debate to improve reasoning and factual accuracy of results which is essential for reliable operations. LLMs have gained prominence as coding assistants [14] demonstrating the capability to function as coders [21], problem solvers [22] or even data scientists [16]. Automated Design of Agentic Systems [13] aims to automatically create AI agent designs by having a meta-agent iteratively program new agents in code.
@@ 3 Features of the System
The third key component is agent autonomy implementable via rules, decision trees, or language models. LLMs are currently preferred due to their natural language interface and intelligent decision-making. LLM high computational needs may cause delays, limiting use in real-time control. The choice of autonomy implementation thus requires balancing LLMs’ flexibility against performance needs in time-sensitive applications.
@@ 4 Examples
In this section, we present two examples, where the aforementioned features can be directly applicable by integrating them into the control systems like [12, 4, 20] and enhancing operational efficiency and stability: the Advanced Light Source (ALS) orbit feedback system and the European XFEL longitudinal feedback manager.
@@ 4.1 ALS Example: Orbit Feedback
Figure 2: Diagram of the orbit feedback agent. The Orbit FB Agent detects abnormal behavior, consults the Logbook Agent for relevant events, and suggests action for the control room to review.
4.2 European XFEL Longitudinal Feedback Manager
The particle beam in a linear accelerator, like the European XFEL, is stabilized within well-defined, narrow parameter spaces through the use of various feedback loops. To effectively orchestrate the complex interplay of these feedback loops, expert systems are often employed [15, 5], which encapsulate the logic of this interaction in software. A Feedback Agent can be trained to learn the possible and desired states, illustrated in Fig. 3. In the next step, such an agent would initially be able to assist the operator as a kind of recommender system and, potentially, could take over the entire operation of all longitudinal feedback loops in a later step.
Figure 3: Diagram of the Longitudinal feedback manager. Left: The operator sets feedback loops and the FB Manager Agent and Logbook Agent observe and recommend actions to the operator. Right: FB Manager Agent takes over the operation and just informs the operator.
@@ 5 Conclusion
This paper proposes a paradigm shift in particle accelerator control through a decentralized multi-agent framework powered by LLMs. By integrating AI agents for high-level tasks and specialized agents for component management, we address the increasing complexity of modern accelerator systems.
@@
Two examples demonstrate the potential of AI agents in assisting complex accelerator tasks, opening the way for higher autonomy. While this approach offers exciting possibilities for improved performance, our practical examples support a vision for more intelligent and adaptive accelerator operations.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction accurately reflect the paper’s contributions and scope. They present the main idea of a decentralized multi-agent framework for accelerator control powered by Large Language Models (LLMs). The introduction outlines the challenges in current accelerator control systems and proposes the paradigm shift, which aligns with the paper’s overall content and examples provided.
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: It primarily focuses on presenting a conceptual framework and discussing potential applications of LLM-powered agents in particle accelerator control, but briefly mentions potentially main issue with LLM latency.
  9. 3.
    Theory Assumptions and Proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [N/A]
  12. Justification: The paper does not include formal theoretical results or proofs. It primarily focuses on presenting a conceptual framework and discussing potential applications of LLM-powered agents in particle accelerator control.
  13. 4.
    Experimental Result Reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  15. Answer: [No]
  16. Justification: The paper does not present experimental results. Instead, it proposes a conceptual framework and provides examples of how it could be applied in specific accelerator control scenarios. As such, there are no experimental results to reproduce.
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  19. Answer: [No]
  20. Justification: The paper does not mention any experiments requiring code or data. It is a conceptual paper proposing a framework for future implementation rather than presenting implemented systems or experiments.
  21. 6.
    Experimental Setting/Details
  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  23. Answer: [No]
  24. Justification: The paper does not include experiments. It presents a conceptual framework and potential applications without conducting specific experiments or providing experimental details.
  25. 7.
    Experiment Statistical Significance
  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  27. Answer: [No]
  28. Justification: The paper does not include experiments or statistical analyses. It is a conceptual paper proposing a framework for future implementation.
  29. 8.
    Experiments Compute Resources
  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  31. Answer: [No]
  32. Justification: The paper does not include experiments or discuss compute resources needed for implementations. It is a conceptual paper proposing a framework for future work.
  33. 9.
    Code Of Ethics
  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  35. Answer: [Yes]
  36. Justification: The paper discusses potential positive and negative societal impacts, which is in line with ethical considerations in AI research.
  37. 10.
    Broader Impacts
  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  39. Answer: [Yes]
  40. Justification: It mentions the potential for improved accelerator performance and efficiency, while also acknowledging potential risks such as the need for human oversight and the importance of safety considerations in autonomous systems.
  41. 11.
    Safeguards
  42. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  43. Answer: [No]
  44. Justification: The paper does not release any data or models that have a high risk for misuse. It is a conceptual paper proposing a framework for future implementation rather than releasing actual systems or data.
  45. 12.
    Licenses for existing assets
  46. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  47. Answer: [Yes]
  48. Justification: The paper cites and credits various existing works and frameworks throughout the text, particularly in the Related Work section. While specific license information is not provided for each reference, the proper attribution of ideas and previous work is evident.
  49. 13.
    New Assets
  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  51. Answer: [No]
  52. Justification: The paper does not introduce or release new assets such as datasets, code, or models. It proposes a conceptual framework for future implementation.
  53. 14.
    Crowdsourcing and Research with Human Subjects
  54. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  55. Answer: [No]
  56. Justification: The paper does not involve crowdsourcing or research with human subjects. It focuses on proposing a conceptual framework for AI-powered accelerator control systems.
  57. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  58. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  59. Answer: [No]
  60. Justification: The paper does not involve research with human subjects, so IRB approval or equivalent is not applicable to this work.

### v4

@@ Abstract
As particle accelerators grow in complexity, traditional control methods face increasing challenges in achieving optimal performance. This paper envisions a paradigm shift: a decentralized multi-agent framework for accelerator control, powered by Large Language Models (LLMs) and distributed among autonomous agents. We present a proposition of a self-improving decentralized system where intelligent agents handle high-level tasks and communication and each agent is specialized to control individual accelerator components.
This approach raises some questions: What are the future applications of AI in particle accelerators? How can we implement an autonomous complex system such as a particle accelerator where agents gradually improve through experience and human feedback? What are the implications of integrating a human-in-the-loop component for labeling operational data and providing expert guidance? We show three examples, where we demonstrate the viability of such architecture.
@@ 1 Introduction
Particle accelerator operation involves a complex interplay of interrelated systems, each requiring precise control and optimization. Traditionally, these systems have been managed through a combination of human expertise and highly specialized algorithms with recent developments including machine learning techniques [8, 9] where some are successfully deployed in operation like e.g[10, 14, 34].
Most of these algorithms are designed to operate in the narrow domain of accelerator functionality. The systems have noise, tight operational limits, and deterministic interplay through the control system to ensure human and equipment safety, preventing scenarios that could lead to hazardous conditions or equipment damage. While effective, this approach can sometimes struggle to achieve optimal overall performance due to the challenges of integrating these disparate systems. Even when a sufficiently comprehensive model of the facility is constructed, it can be susceptible and prone to problems as the accelerator transitions through different states or experiences drifts and thus requires human intervention for setup and ongoing maintenance.
This paper proposes a paradigm shift in accelerator control systems, envisioning a future where self-improving agents control separated sub-components using algorithms that inform operators of significant events. In this framework, we show how autonomous agents controlled by LLMs serve not only in the creation and retrieval of accelerator operation documentation but also as high-level coordinators, handling complex tasks such as interpreting the machine state from analyzing control system channels or inter-agent communication with a degree of autonomy in decision-making and communication. Meanwhile, dedicated agents controlled by LLMs execute specific, well-established calculations or operations within their domains of expertise. This decentralized approach, illustrated in our conceptual diagram Fig. 1, showcases a potential future for the operation of particle accelerators with the flexibility to deploy more advanced agents.
@@
Figure 1: Schematic overview of a decentralized agent-based control architecture. The left diagram illustrates the modular structure where each agent, consisting of devices, algorithms, and agent controllers, manages specific subsystems. The right figure shows these agents interacting with both the physical components of the accelerator and the control room, enabling prompting current state, monitoring, decision-making, and adaptive responses to complex operational scenarios.
@@ 2 Related Work
Machine learning techniques have increasingly been applied to various aspects of accelerator physics in recent years. Edelen et al[8, 9] provide a comprehensive review of machine learning applications in particle accelerators, covering areas such as beam diagnostics, control systems, and performance optimization.
@@ 2.1 Agentic AI and LLMs in Complex Reasoning Tasks
There is a rise of agentic AI that has demonstrated remarkable capabilities in complex reasoning tasks [3]. According to [31], an effective agent in an agentic system should possess autonomy in independent operation, reactivity in environmental response, and proactiveness in pursuing its objectives, enabling it to function intelligently and adaptively in complex environments. We will highlight the most relevant works that align with our problem.
@@
The use of AI agents to effectively engage with their environment and complete a wide array of tasks has been gaining increasing popularity. Park et al[22] introduce generative agents that use LLMs to simulate human behavior in interactive environments, showcasing how LLMs can be utilized to model intricate multi-agent interactions and environmental complexities. Wang et al[28] uses an LLM-powered agent to autonomously explore, acquire skills, and make discoveries in unfamiliar environments using components like an automatic curriculum and skill library. AgentVerse [4] shows a framework for collaborative groups of LLM-powered expert agents, showing improved performance over single agents in various tasks. This multi-agent approach leverages LLMs for high-level tasks while enabling specialized components to work together autonomously on complex problems. AutoGen [30] framework allows building LLM applications using multiple conversable agents. They show that autonomous, communicating agents can control subcomponents of a system. Particle accelerator control often utilizes operation sequencing [2, 11], there LLM+P [20] can help by translating natural language into formal planning languages, potentially bridging the gap between planning, natural language, and accelerator controls. To deal with environment changes, Shinn et al[26] proposes Reflexion, a framework that reinforces language agents through linguistic feedback rather than traditional weight updates. Reflexion agents verbally reflect on task feedback and maintain reflective text in an episodic memory buffer to improve decision-making in subsequent trials. To impose the factuality, [7] proposes multiagent debate where multiple LLMs debate to improve reasoning and factual accuracy of results which is essential for reliable operations. LLMs have gained prominence as coding assistants [17] demonstrating the capability to function as coders [24], problem solvers [25] or even data scientists [19]. Automated Design of Agentic Systems [16] aims to automatically create AI agent designs by having a meta-agent iteratively program new agents in code.
@@ 3 Features of the System
The third key component is agent autonomy implementable via rules, decision trees, or language models. LLMs are currently preferred due to their natural language interface and intelligent decision-making. LLM’s high computational needs may cause delays, limiting use in real-time control. The choice of autonomy implementation thus requires balancing LLMs’ flexibility against performance needs in time-sensitive applications.
@@ 4 Examples
In this section, we present three examples, where the aforementioned features can be directly applicable by integrating them into the control systems like [5, 15, 23] and enhancing operational efficiency and stability: the Advanced Light Source (ALS) orbit feedback system, the European XFEL longitudinal feedback manager and Fermi coding assistant.
@@ 4.1 ALS Example: Orbit Feedback
Figure 2: Diagram of the orbit feedback agent. The Orbit FB Agent detects abnormal behavior, consults the Logbook Agent for relevant events, and suggests an action for the control room to review.
4.2 European XFEL Example: Longitudinal Feedback Manager
The particle beam in the linear accelerator, like the European XFEL, is stabilized within well-defined, narrow parameter spaces through the use of various feedback loops. To effectively orchestrate the complex interplay of these feedback loops, expert systems are often employed [18, 6], which encapsulate the logic of this interaction in software. A Feedback Agent can be trained to learn the possible and desired states, illustrated in Fig. 3. In the next step, such an agent would initially be able to assist the operator as a kind of recommender system and, potentially, could take over the entire operation of all longitudinal feedback loops in a later step.
Figure 3: Diagram of the Longitudinal feedback manager. Left: The operator sets feedback loops and the FB Manager Agent and Logbook Agent observe and recommend actions to the operator. Right: FB Manager Agent takes over the operation and informs the operator.
@@ 4.2 European XFEL Example: Longitudinal Feedback Manager
4.3 Fermilab Example: ACL Coding Assistant
Accelerator Command Language (ACL), developed at Fermilab, is a scripting language designed for non-programmers, such as engineers and accelerator operators, to control complex accelerator systems. Despite its specialized nature, AI agents can assist with ACL, even without specific training on the language itself.
A Code Retrieval Agent utilizes a preexisting semantic search API to query ACL’s extensive documentation, retrieving relevant code snippets based on detailed descriptions. After gathering these results, the Retrieval Agent reviews the information and selects the most relevant snippets to fulfill the user’s request. These snippets are then passed to a Code Generation Agent, which uses the retrieved information to infer the correct ACL code structure and generate context-aware script suggestions using e.g[12].
This two-step process simplifies working with ACL and provides intelligent coding assistance, helping users efficiently generate scripts for a proprietary language specific to Fermilab.
However, one potential challenge is that the initially generated code may not always be correct. To address this, combining Reflexion [26] for verbal reinforcement learning and embeddings for long-term memory such as [33] could help guide the model toward generating fully functional code. This approach would enable the model to iteratively refine its outputs based on feedback and past experiences, ensuring more accurate results over time.
5 Safety Issues and Hallucinations
The tendency of LLMs to hallucinate creates substantial risks for critical system applications. To mitigate potential issues arising from hallucinations and uncertainty, several approaches can be implemented. These include constraining output using defined grammars or regular expressions [29], employing multiple agents to cross-check decisions [1], and integrating real-time sensor data feedback loops.
It’s important to emphasize that this proposal is for an AI-based, high-level control system that would operate alongside existing control systems. These existing systems have built-in safety measures for protecting equipment and personnel, which would remain independent of the proposed AI control. Our approach is to begin with passive monitoring and operator suggestions before progressing to limited control of non-critical systems.
Furthermore, fallback mechanisms should automatically revert to traditional control systems if anomalies are detected. Continuous validation against established algorithms and expert decisions will be an integral part of the proof of concept. The primary role of the system in this context remains sensing and informing human operators rather than autonomous control, thereby maintaining critical human oversight in accelerator operations.
@@ 6 Conclusion
This paper shows a paradigm shift in particle accelerator control through a decentralized multi-agent framework powered by LLMs. By integrating AI agents for high-level tasks and specialized agents for component management, we address the increasing complexity of modern accelerator systems.
@@
Three examples demonstrate the potential of AI agents in assisting complex accelerator tasks, opening the way for higher autonomy and posing a question of the use case of intelligent agents for assistance with operating particle accelerators. This approach offers exciting possibilities for improved performance and our practical examples showcase a vision for more intelligent and adaptive accelerator operations.